OpenStudy (anonymous):

Exponetial Decay is killing me plz help A 33 gram sample of a substance that's used for drug research has a k-value of 0.1081. N = N0e -kt N0 = initial mass (at time t = 0) N = mass at time t k = a positive constant that depends on the substance itself and on the units used to measure time t = time, in days Find the substance's half-life, in days. Round your answer to the nearest tenth.

OpenStudy (anonymous):

The half life is when: \(N_0 /2 = N_0 e^{-kt}\) \(e^{-kt} = 1/2\) \(-kt = ln(1/2) = -ln(2)\) \(t = \frac{ln(2)}{k}\)

OpenStudy (anonymous):

\(ln(2) \approx 0.693\)